OpenStudy (anonymous):

The Pentagon is a pentagonal-shaped building near Washington, D.C., that is home to the United States Department of Defense. Write a formula for the perimeter of a regular pentagon in terms of the length of a side. Evaluate the formula to find the perimeter of the Pentagon, which has a side length of 921 ft. - and what is the side length of a pentagon?

OpenStudy (calculusxy):

perimeter=ns. n; number of sides,s; lenght of side. perimeter=5(921). perimeter=4605ft

OpenStudy (ybarrap):

Perimeter (P) is sum of all sides (x): $$ P=x+x+x+x+x=5\times x\\=5\times 921 $$ 

OpenStudy (anonymous):

so do i multiply 5 x 921? which is 4605

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

yes maam

OpenStudy (anonymous):

so the side length is 4605? :) @ybarrap

OpenStudy (ybarrap):

4605 is the perimeter and the length of one side is 921.

OpenStudy (anonymous):

ok thankyou!

OpenStudy (ybarrap):

yw